OpenStudy (fluttershyk):

Find the Circumference of the circle (use 3.14 for pi). Show your work. Round to the nearest tenth. @lyss @Will.H

OpenStudy (will.h):

C = 2 * pi * r C = 2 * 3.14 * r what is r?

OpenStudy (lyss):

Circumference = pi x diameter C = pi x radius x 2 C = 3.14 x radius of circle in image x 2

OpenStudy (will.h):

i think you forgot to post the figure

OpenStudy (fluttershyk):

Sorry it was the same one as last time

OpenStudy (fluttershyk):

@lyss and @Will.H

OpenStudy (will.h):

C = 2 * 3.14 * r C = 2 * 3.14 * 23 C = 144.4

OpenStudy (fluttershyk):

Thank you

OpenStudy (will.h):

you welcome